OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x − 7 y = −one halfx − 2 y = 2x − 2 y = −one halfx − 7

OpenStudy (anonymous):

ok so if it's parallel the slope has to be the same, so 2x

OpenStudy (anonymous):

next you know a point so you plug in the slope and the point given into y=mx+b

OpenStudy (anonymous):

sooo..... 3=2(5)+b 3=10+b -10 -10 -7=b

OpenStudy (anonymous):

y=2x-7

OpenStudy (anonymous):

any qs?